Let X1,X2,...,Xn be independent and identifically distributed random variables each having pdf



f(x)=q^xP .x=0,1,2....



P+q=1



Let Sn=X1+X2+...+Xn.



Obtain the pgf of Sn and find the mean and variance of Sn



1
Expert's answer
2022-01-10T13:32:06-0500

Solution:

X1,X2,...,Xn be IID.

"f(x)=q^xp;,x=0,1,2,...
\\X\sim G(p)
\\E(X)=\dfrac qp, Var(X)=\dfrac q{p^2}
\\ M_X(t)=\dfrac p{1-qe^t}, G_X(t)=\dfrac p{1-qt}
\\S_n=X_1+X_2+...+X_n
\\ G_{S_n}(Z)=G_{X_1}(Z)G_{X_2}(Z)...G_{X_n}(Z)=\Pi_{i=1}^nG_{X_i}(Z)
\\G_X(Z)=E(Z^x)=\dfrac p{1-qz}\ \ \ \ \ \ [G_X(.)\rightarrow pdf\ of\ X]"

So, "G_{S_n}(Z)=(\dfrac p{1-qz})^n"

"E(S_n)=E(X_1+X_2+...+X_n)
\\=\Sigma_{i=1}^nE(X_i)
\\=\Sigma_{i=1}^n \dfrac qp
\\=\dfrac {nq}p"

"Var(S_n)=Var(X_1+X_2+...+X_n)
\\=\Sigma_{i=1}^nVar(X_i)" [Since Xi's are IIDs]

"=n\times \dfrac q{p^2}
\\=\dfrac {nq}{p^2}"